OpenStudy (anonymous):

Write each equation in slope intercept form: 3x-5y=5

OpenStudy (anonymous):

x int let y=0 3x=5 x= 5/3 y-int let x =0 -5y=5 y=-1

OpenStudy (anonymous):

5y = 3x-5 y = (3/5)x - 1

OpenStudy (anonymous):

Thank you